OpenStudy (dominirican1013):

Evaluate the following expression: -1(5,-1) + 2(1,1) A. (5,-2) B. (-3,3) C. (5,-2) D. (7,1)

OpenStudy (mathmate):

Hint: use the scalar multiplication property: s(a,b)=(sa+sb). For example, 5(-2,1)=(5(-2),5*1)=(-10,5)

OpenStudy (dominirican1013):

I need help with a different question

OpenStudy (anonymous):

kool what is it

OpenStudy (dominirican1013):

What is the length (magnitude) of the vector (-3,2) A. \[\sqrt{2}\] B. \[\sqrt{13}\] C. 0 D. 3

OpenStudy (dominirican1013):

Does anyone know this?

OpenStudy (anonymous):

im working it out hold up please

OpenStudy (dominirican1013):

it says im typing but i wasnt lol

OpenStudy (dominirican1013):

is it square root of 13

OpenStudy (dominirican1013):

It was square root of 13 thank you for trying to help though :D

OpenStudy (anonymous):

ight finally its and nvm u got it i tried damn

OpenStudy (dominirican1013):

You gt a medal for trying though :D

OpenStudy (dominirican1013):

I have another question though

OpenStudy (dominirican1013):

Evaluate the dot product of (2,4) and (1,-2)